OpenStudy (anonymous):

The ratio of foreign students to American students at a college is 2 to 35. How many foreign students attend this college if there are 1575 American students?

jimthompson5910 (jim_thompson5910):

let x = number of foreign students

jimthompson5910 (jim_thompson5910):

since "The ratio of foreign students to American students at a college is 2 to 35." and because " there are 1575 American students" we can say this ratio = (number of foreign students)/(number of American students) 2/35 = x/1575 we then cross multiply to get 2*1575 = 35*x I'll let you finish up

OpenStudy (anonymous):

i dont get

jimthompson5910 (jim_thompson5910):

you want to isolate x so how would you do that

OpenStudy (anonymous):

90

jimthompson5910 (jim_thompson5910):

correct

jimthompson5910 (jim_thompson5910):

since we let x = number of foreign students and because x = 90, we now know there are 90 foreign students